Find the slope of the curve y=11x^2 at (-4,176)
Answer by ikleyn(52777)

.
Find the slope of the curve y=11x^2 at (-4,176)
~~~~~~~~~~~~~~~~~~~~

The slope of the curve  y=11x^2  at (-4,176) is equal to the value of the derivative of the function at  x = -4.


The derivative is  y'(x) = 2*11*x = 22*x.


The value of the derivative at x = -4 is  y'(-4) = 22*(-4).


Calculate it on your own.


This value is your answer.